# Quantitative description of strongly correlated materials by combining downfolding techniques and tensor networks

### v1

@@ Abstract
We present a high-accuracy procedure for electronic structure calculations of strongly correlated materials. The approach involves downfolding the full Hilbert space into a low-energy subspace that captures the most significant electron correlations, leading to an effective Hubbard model. This generalized model is then solved using tensor network methods. Our work focuses on one-dimensional and quasi-one-dimensional materials, for which we employ the machinery of matrix product states. We apply this framework to the conjugated polymers trans-polyacetylene and polythiophene, as well as the quasi-one-dimensional charge-transfer insulator Sr2CuO3. The predicted band gaps exhibit quantitative agreement with state-of-the-art computational techniques and experimental measurements. Beyond band gaps, tensor networks provide access to a wide range of physically relevant properties, including spin magnetization and various excitation energies. Their flexibility supports the implementation of complex Hamiltonians with longer-range interactions, while the bond dimension enables systematic control over accuracy. Furthermore, the computational cost scales efficiently with system size, demonstrating the framework’s scalability.
@@ I Introduction
Simulating molecules and materials from first principles is essential not only for predicting properties of interest but also for understanding the mechanisms behind emergent phenomena. Techniques like density functional theory (DFT) [1, 2] have enabled the quantitative description of the electronic structure of many materials [3, 4]. The local density approximation (LDA) and generalized gradient approximation (GGA) to the DFT exchange-correlation functional are commonly used to investigate weakly correlated systems. However, the accuracy of these estimations is often insufficient [5, 6], particularly for strongly correlated materials, which exhibit some of the most intriguing and least understood features, such as unconventional superconductivity [7, 8, 9, 10, 11, 12], colossal magnetoresistance [13, 14, 15], and metal-insulator transitions [16, 17].
@@
Alternatively, downfolding methods aim to construct an effective model that captures the essential physics by retaining only the low-energy degrees of freedom [25]. A suitable target space typically includes states that dominate the material’s behavior at the low-energy scale. The resulting effective model, often represented as a Hubbard model [32], can then be solved with greater accuracy to extract key physical properties. Although the concept is straightforward, several technical challenges complicate the procedure. One key issue, common to most downfolding routines, is double counting [33]. The screening effects by electrons outside the target space should be incorporated into the two-body interaction parameters, while screening within the low-energy space must be excluded, as it is already accounted for in the effective model. Various constrained methods have been developed to deal with this issue [34, 35, 36, 37, 38, 39, 40]. Among them, the constrained random phase approximation (cRPA) [41, 42] is perhaps the most widely used technique for calculating the screened interaction parameters in solids, in part due to its ability to retrieve the frequency-dependence [43]. A second source of double counting arises from the fact that one-body hopping terms are determined from an initial computation that already includes some electron interactions through the Hartree term and the exchange-correlation functional. To address this, correction terms can be subtracted from the parameters. Exclusion of the Hartree contribution within the target space is relatively straightforward, but correcting for the exchange-correlation is more difficult. While multiple correction schemes exist [20, 44, 45, 46], their impact on the accuracy of simulations remains questionable [33, 47, 48]. On the other hand, a constrained (c) calculation effectively removes double counting up to the perturbation order within the scheme [39].
@@
Once the model is constructed, the next challenge is solving it. Substantial progress has been made with dynamical mean field theory (DMFT) [47, 49, 50], which maps the Hubbard model to an Anderson impurity model [51]. Impurity models can be solved with even greater precision using methods like quantum Monte Carlo (QMC) [52, 53] or coupled-cluster techniques. Conventional DFT+DMFT [47, 20, 54, 55] has been modified to extended DMFT (EDMFT), +DMFT, and even +EDMFT, offering solutions to some of the initial limitations [56, 57, 58, 59]. However, challenges remain, such as the absence of non-local spin fluctuations and non-local vertex corrections beyond RPA [25]. Ongoing research is exploring further extensions to overcome these issues [60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70]. Besides DMFT, other methods like density matrix embedding theory (DMET) [71] and self-energy embedding theory (SEET) [72, 73], which are closely related to DMFT, have been developed, alongside the direct application of solvers such as QMC to the constructed Hubbard model. Directly solving the Hubbard model allows for a more intuitive interpretation as it closely resembles the physical system. Furthermore, the approximations introduced by DMFT and DMET can only be improved by increasing the size of the cluster, leading to exponential growth of the computational cost.
@@
While the applicability of QMC is limited by the sign problem [74], coupled cluster methods are generally impractical for realistic multi-band Hubbard models beyond the singles and doubles (CCSD) approximation. In the past two decades, tensor networks (TNs) [75] have emerged as powerful tools for describing interacting quantum lattice systems. They are applicable to generalized Hubbard models with longer-range interactions while maintaining a computational cost that scales favorably with system size according to the area law of entanglement [76, 75]. Additionally, the bond dimension is a tunable parameter for simulation accuracy, with infinite bond dimension representing the exact solution, though convergence is often achievable with relatively small, feasible values. Serious attention was first drawn to TNs after the success of the density matrix renormalization group (DMRG) [77] algorithm, now well established [78]. DMRG efficiently optimizes matrix product states (MPS) [79, 80, 81], a one-dimensional TN Ansatz. Algorithms based on projected entangled-pair states (PEPS) [82, 76] show increasing promise for higher-dimensional systems, particularly for fermionic systems where the sign problem excludes QMC. However, these higher-dimensional TN methods remain relatively new compared to many other techniques, and intensive research continues to broaden their performance and applicability.
@@
In this work, we use DFT+cRPA to construct an effective multi-band Hubbard model for several one-dimensional (1D) and quasi-1D materials, which we then solve using MPS methods. Conjugated polymers provide ideal benchmark systems for our approach [83]. These organic semiconductors were extensively investigated during the late 20th century [84, 85, 86, 87, 88, 89], leading to the Nobel Prize in Chemistry in 2000 for the discovery and development of conductive polymers. With their continued relevance in modern applications such as flexible electronics, biosensors, and light-emitting diodes [83, 90], conjugated polymers remain of significant interest. The band gap, obtainable via TN calculations, will be an important quantity to validate the procedure. Although optical and band gap measurements are often limited to bulk samples, angle-resolved photoemission spectroscopy (ARPES) experiments on individual chains have recently been demonstrated [91]. Specifically, we will study trans-polyacetylene (tPA) and polythiophene (PT), due to their extensive presence in the literature. Their crystal structures are shown in Fig. 1a and b.
@@
In addition to purely 1D materials, we examine the quasi-1D system Sr2CuO3. As illustrated in Fig. 1c, infinite chains of corner-sharing CuO4 squares are embedded within its three-dimensional crystal structure [92]. These chains are well-isolated, allowing the low-energy physics to be effectively treated as one-dimensional. Sr2CuO3 may allow insights in the mechanisms behind superconducting cuprates through oxygen doping [93]. It also provides a rich platform for exploring spin-charge separation, where single-particle excitations split into spinons and holons carrying spin and charge, respectively [94, 95], as observed in ARPES experiments [96, 97]. Its antiferromagnetic spin structure along the chain and the elementary excitations are observables that are accessible through TN calculations of the Hubbard model.
@@
The purpose of this paper is twofold. First, we demonstrate the effectiveness of tensor networks in modeling real materials. Although our focus is on one-dimensional systems, the transition to higher-dimensional PEPS calculations, anticipated in the near future, is expected to follow a similar framework. Second, we introduce a novel ab initio simulation technique for strongly correlated electron systems, applicable to both purely 1D materials and those exhibiting a strong one-dimensional character. This approach not only captures electron correlation accurately through effective models with a high number of bands and longer-range interactions, but also provides access to various excitations. Additionally, the bond dimension offers systematic control over accuracy, with computational costs scaling efficiently as system size increases. The next section outlines our approach. Section III presents the results, while we draw our conclusions in Section IV. Finally, Section V elaborates on the computational details.
@@ II Methodology
The central problem of the current work is the many-body problem, and more specifically, the electronic structure of strongly correlated materials. The proposed procedure consists of two main components: constructing the Hubbard model and solving it. Fig. 2 illustrates the various steps of our method. We begin by explaining the theory behind the downfolding algorithm, followed by a description of how tensor networks are applied to solve the low-energy model.
@@ II.1.1 DFT and target space
The starting point of the procedure is an initial DFT calculation of the material. With a relatively cheap exchange-correlation functional, like the Perdew-Burke-Ernzerhof (PBE)-GGA [99], the Kohn-Sham orbitals and corresponding eigenvalues can be calculated, leading to an approximate band structure. Assuming that DFT provides sufficient accuracy when correlations are weak, a subset of bands should be chosen to define the low-energy space where strong correlation effects are incorporated in the effective model. In order to keep the computational cost low, this target space is preferably as small as possible without neglecting relevant physics. Typically, correlations are expected to be strong in narrow and partially occupied bands close to the Fermi energy. For bands further away from the Fermi energy, it is assumed that correlations are weak and, consequently, that DFT is accurate.
@@
is a linear combination of Bloch functions , which can be approximated by the Kohn-Sham orbitals. For a spin-polarized DFT calculation, the spin is included in the orbital index. When the Brillouin zone is discretized into a set of -points, as is the case in DFT calculations, Equation (2) becomes
@@ II.1.2 Coulomb integrals
The main idea of the cRPA is then to split this sum into the transitions within the chosen low-energy subspace and the remaining transitions, . As the screening due to states within the target space is included in the effective model, we exclude them from the screened interaction to arrive at a partially screened interaction,
@@ II.1.3 Hopping parameters
with the Kohn-Sham Hamiltonian. However, the Hartree term and exchange-correlation energy should be constrained to exclude the interactions within the target space. The low-energy Hartree contribution to parameter is given by [107, 108, 45]
with the partially screened Hartree term and the single-particle occupation matrix in the MLWF basis, obtained from DFT. One can choose to sum only over the interaction parameters included in the effective model, treating the remaining interactions in a mean-field way, or to sum over all . In the latter case, the sum is truncated for large and , since the interaction parameters decay with distance. We employ this approach, exclusively incorporating interactions into the two-body terms.
@@ II.2.1 Matrix product states
Any quantum state of a lattice model with sites can be expressed as an expansion over basis states,
@@
The column vectors and handle the boundary conditions, but carry no physical meaning as we work in an infinite system. In Equation (18), we introduced the diagrammatic notation for tensor networks. Shapes represent tensors and legs their indices. Here, the legs pointing downward correspond to the physical Hilbert space of dimension , while the horizontal legs denote the virtual spaces of dimension . Connected legs imply contraction of the indices. Generally, the unit cell can be expanded to multiple sites, forming a similar chain but with different repeating tensors.
@@ II.2.3 Quasiparticle excitations
In addition to the system’s ground state, we are interested in elementary excitations, which enable the calculation of the band gap. The quasiparticle Ansatz in the context of MPS, as introduced by Haegeman et al. [115, 116], is given by the expression
@@
with a vectorized version of the matrix . The effective Hamiltonian and normalization matrix are defined by the relations
@@
III Results and discussion
We present the results of our approach applied to tPA, PT, and Sr2CuO3. For tPA, a relatively simple system, we focus on the flexibility and quantitative accuracy that is offered by TNs. In the case of PT, we demonstrate the feasibility of effective models containing a large number of bands. Finally, we show that our approach is also applicable to quasi-1D materials such as Sr2CuO3. More information on the computational details can be found in Section V.
III.1 Trans-polyacetylene
Polyacetylene is the simplest conjugated polymer, existing in two configurations: trans- (see Fig. 1a) and cis-polyacetylene. The cis isomer is thermodynamically less stable and converts to the trans form upon heating, allowing the preparation of pure trans samples for experimental studies [118]. Therefore, our discussion focuses on trans-polyacetylene.

### v2

@@ Abstract
We present a high-accuracy procedure for electronic structure calculations of strongly correlated materials. To address limitations in current electronic structure methods, we employ density functional theory in combination with the constrained random phase approximation to construct an effective multi-band Hubbard model, which is subsequently solved using tensor networks. Our work focuses on one-dimensional and quasi-one-dimensional materials, for which we employ the machinery of matrix product states. We apply this framework to the conjugated polymers trans-polyacetylene and polythiophene, as well as the quasi-one-dimensional charge-transfer insulator Sr2CuO3. The predicted band gaps show quantitative agreement with state-of-the-art computational techniques and experimental measurements. Beyond band gaps, tensor networks provide access to a wide range of physically relevant properties, including spin magnetization and various excitation energies. Their flexibility supports the implementation of complex Hamiltonians with longer-range interactions, while the bond dimension enables systematic control over accuracy. Furthermore, the computational cost scales efficiently with system size, demonstrating the framework’s scalability.
@@ I Introduction
Simulating molecules and materials from first principles is essential not only for predicting properties of interest but also for understanding the mechanisms behind emergent phenomena. Techniques like density functional theory (DFT) [1, 2] have enabled the quantitative description of the electronic structure of many materials [3, 4]. The local density approximation (LDA) and generalized gradient approximation (GGA) to the DFT exchange-correlation functional are commonly used to investigate weakly correlated systems. However, the accuracy of these estimations is often insufficient [5, 6], particularly for strongly correlated materials, which exhibit some of the most intriguing and least understood features, such as unconventional superconductivity [7, 8, 9, 10, 11], colossal magnetoresistance [12, 13, 14], and metal-insulator transitions [15, 16].
@@
Alternatively, downfolding methods aim to construct an effective model that captures the essential physics by retaining only the low-energy degrees of freedom, typically the electronic states close to the Fermi level [24]. A suitable target space includes these states, which dominate the material’s behavior at the low-energy scale. The resulting effective model, often represented as a Hubbard model [31], can then be solved with greater accuracy to extract key physical properties. Although the concept is straightforward, several technical challenges complicate the procedure. One key issue, common to most downfolding routines, is double counting [32]. The screening effects by electrons outside the target space should be incorporated into the two-body interaction parameters, while screening within the low-energy space must be excluded, as it is already accounted for in the effective model. Various constrained methods have been developed to deal with this issue [33, 34, 35, 36, 37, 38, 39]. Among them, the constrained random phase approximation (cRPA) [40, 41] is perhaps the most widely used technique for calculating the screened interaction parameters in solids, in part due to its ability to retrieve the frequency-dependence [42]. A second source of double counting arises from the fact that one-body hopping terms are determined from an initial computation that already includes some electron interactions through the Hartree term and the exchange-correlation functional. To address this, correction terms can be subtracted from the parameters. Exclusion of the Hartree contribution within the target space is relatively straightforward, but correcting for the exchange-correlation is more difficult. While multiple correction schemes exist [19, 43, 44, 45], their impact on the accuracy of simulations remains questionable [32, 46, 47]. On the other hand, a constrained (c) calculation effectively removes double counting up to the perturbation order within the scheme [38].
@@
Once the model is constructed, the next challenge is solving it. Substantial progress has been made with dynamical mean field theory (DMFT) [46, 48, 49], which maps the Hubbard model to an Anderson impurity model [50]. Impurity models can be solved with even greater precision using methods like quantum Monte Carlo (QMC) [51, 52] or coupled-cluster techniques. Conventional DFT+DMFT [46, 19, 53, 54] has been modified to extended DMFT (EDMFT), +DMFT, and even +EDMFT, offering solutions to some of the initial limitations [55, 56, 57, 58]. However, challenges remain, such as the absence of non-local spin fluctuations and non-local vertex corrections beyond RPA [24]. Ongoing research is exploring further extensions to overcome these issues [59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69]. Besides DMFT, other methods like density matrix embedding theory (DMET) [70, 71] and self-energy embedding theory (SEET) [72, 73], which are closely related to DMFT, have been developed, alongside the direct application of solvers such as QMC to the constructed Hubbard model. Directly solving the Hubbard model allows for a more intuitive interpretation as it closely resembles the physical system. A key limitation of DMFT and DMET is that improving their approximations primarily relies on increasing cluster size, leading to exponential growth of the computational cost. Recent work has focused on systematic improvability in quantum embedding that extends beyond merely enlarging fragment sizes [74].
@@
While the applicability of QMC is limited by the sign problem [75], coupled cluster methods are generally impractical for realistic multi-band Hubbard models beyond the singles and doubles (CCSD) approximation. In the past two decades, tensor networks (TNs) [76] have emerged as powerful tools for describing interacting quantum lattice and molecular systems [77, 78, 79, 80, 81, 82]. They are applicable to generalized Hubbard models with longer-range interactions while maintaining a computational cost that scales favorably with system size according to the area law of entanglement [83, 76]. Additionally, the bond dimension is a tunable parameter for simulation accuracy, with infinite bond dimension representing the exact solution, though convergence is often achievable with relatively small, feasible values. Serious attention was first drawn to TNs after the success of the density matrix renormalization group (DMRG) [84] algorithm, now well established [85]. DMRG efficiently optimizes matrix product states (MPS) [86, 87, 88], a one-dimensional TN Ansatz. Algorithms based on projected entangled-pair states (PEPS) [89, 83] show increasing promise for higher-dimensional systems, particularly for fermionic systems where the sign problem excludes QMC. However, these higher-dimensional TN methods remain relatively new compared to many other techniques, and intensive research continues to broaden their performance and applicability.
@@
In this work, we use DFT+cRPA to construct an effective multi-band Hubbard model for several one-dimensional (1D) and quasi-1D materials, which we then solve using MPS methods. Conjugated polymers provide ideal benchmark systems for our approach [90]. These organic semiconductors were extensively investigated during the late 20th century [91, 92, 93, 94, 95, 96], leading to the Nobel Prize in Chemistry in 2000 for the discovery and development of conductive polymers. With their continued relevance in modern applications such as flexible electronics, biosensors, and light-emitting diodes [90, 97], conjugated polymers remain of significant interest. The band gap, obtainable via TN calculations, will be an important quantity to validate the procedure. Although optical and band gap measurements are often limited to bulk samples, angle-resolved photoemission spectroscopy (ARPES) experiments on individual chains have recently been demonstrated [98]. Specifically, we will study trans-polyacetylene (tPA) and polythiophene (PT), due to their extensive presence in the literature. Their crystal structures are shown in Figure 1a and b.
@@
In addition to purely 1D materials, we examine the quasi-1D system Sr2CuO3. As illustrated in Figure 1c, infinite chains of corner-sharing CuO4 squares are embedded within its three-dimensional crystal structure [99]. These chains are well-isolated, allowing the low-energy physics to be effectively treated as one-dimensional. Sr2CuO3 may allow insights in the mechanisms behind superconducting cuprates through oxygen doping [100]. It also provides a rich platform for exploring spin-charge separation, where single-particle excitations split into spinons and holons carrying spin and charge, respectively [101, 102], as observed in ARPES experiments [103, 104]. Its antiferromagnetic spin structure along the chain and the elementary excitations are observables that are accessible through TN calculations of the Hubbard model.
@@
The purpose of this paper is twofold. First, we demonstrate the effectiveness of tensor networks in modeling real materials. Although our focus is on one-dimensional systems, the transition to higher-dimensional PEPS calculations, anticipated in the near future, is expected to follow a similar framework. Second, we introduce a novel ab initio simulation technique for strongly correlated electron systems, applicable to both purely 1D materials and those exhibiting a strong one-dimensional character. This approach not only captures electron correlation accurately through effective models with a high number of bands and longer-range interactions, but also provides access to various excitations. Additionally, the bond dimension offers systematic control over accuracy, with computational costs scaling efficiently as system size increases. The next section outlines our approach, followed by the computational details in Section III. Section IV presents the results, while we draw our conclusions in Section V.
@@ II Methodology
The central problem of the current work is the many-body problem, and more specifically, the electronic structure of strongly correlated materials. The proposed procedure consists of two main components: constructing the Hubbard model and solving it. Figure 2 illustrates the various steps of our method. We begin by explaining the theory behind the downfolding algorithm, followed by a description of how tensor networks are applied to solve the low-energy model.
@@ II.1.1 DFT and Target Space
The starting point of the procedure is an initial DFT calculation of the material. With a relatively cheap exchange-correlation functional, like the Perdew-Burke-Ernzerhof (PBE)-GGA [106], the Kohn-Sham orbitals and corresponding eigenvalues can be calculated, leading to an approximate band structure. Assuming that DFT provides sufficient accuracy when correlations are weak, a subset of bands should be chosen to define the low-energy space where strong correlation effects are incorporated in the effective model. In order to keep the computational cost low, this target space is preferably as small as possible without neglecting relevant physics. When feasible, more bands may be included to increase the accuracy. Typically, correlations are expected to be strong in narrow and partially occupied bands close to the Fermi energy. For bands further away from the Fermi energy, it is assumed that correlations are weak and, consequently, that DFT is accurate [19, 24].
@@
is a linear combination of Bloch functions , which can be approximated by the Kohn-Sham orbitals. For a spin-polarized DFT calculation, the spin is included in the orbital index. When the Brillouin zone is discretized into a set of -points, as is the case in DFT calculations, eq (2) becomes
@@ II.1.2 Coulomb Integrals
The main idea of the cRPA is then to split this sum into the transitions within the chosen low-energy subspace and the remaining transitions, . As the screening due to states within the target space is included in the effective model, we exclude them from the screened interaction to arrive at a partially screened interaction
@@ II.1.3 Hopping Parameters
with the Kohn-Sham Hamiltonian. However, the Hartree term and exchange-correlation energy should be constrained to exclude the interactions within the target space. The low-energy Hartree contribution to parameter is given by [114, 115, 116, 44]
with the partially screened Hartree term, the electron density, and the single-particle occupation matrix in the MLWF basis, obtained from DFT. One can choose to sum only over the interaction parameters included in the effective model, treating the remaining interactions in a mean-field way, or to sum over all . In the latter case, the sum is truncated for large and , since the interaction parameters decay with distance. We employ this approach, exclusively incorporating interactions into the two-body terms.
@@ II.2.1 Matrix Product States
Any quantum state of a lattice model with sites can be expressed as an expansion over basis states
@@
The column vectors and handle the boundary conditions, but carry no physical meaning as we work in an infinite system. In eq (18), we introduced the diagrammatic notation for tensor networks. Shapes represent tensors and legs their indices. Here, the legs pointing downward correspond to the physical Hilbert space of dimension , while the horizontal legs denote the virtual spaces of dimension . Connected legs imply contraction of the indices. Generally, the unit cell can be expanded to multiple sites, forming a similar chain but with different repeating tensors.
@@ II.2.3 Quasiparticle Excitations
In addition to the system’s ground state, we are interested in elementary excitations, which enable the calculation of the band gap. The quasiparticle Ansatz in the context of MPS, as introduced in Refs. [123, 124], is given by the expression
@@
with a vectorized version of the matrix and the excitation energy. The effective Hamiltonian and normalization matrix are defined by the relations
@@
III Computational Details
III.1 Structure Relaxation
Geometry relaxation is performed using VASP [127, 128] with the projector augmented-wave method [129]. For tPA, structures labeled tPA1 and tPA2 were obtained using the PBE [106] and B3LYP [20] functionals, respectively, each with a 20 Å vacuum layer to isolate the chains. A plane-wave energy cutoff of 600 eV is applied on a -grid along the chain axis. Electronic minimization proceeds until the energy difference falls below eV, while structural relaxation continues until forces converge to eV.
III.2 Downfolding
Downfolding for tPA and PT is performed via 1D DFT calculations in PySCF [130], employing the PBE functional. Although more advanced functionals could potentially yield a more accurate band structure, it is not a priori clear whether they would produce improved model parameters due to double-counting issues. A systematic in depth investigation would be very interesting but is beyond the study of the current work. For tPA, the all-electron correlation-consistent polarized valence triple-zeta (cc-pVTZ) basis set [131] is used, while for PT, the pseudopotential GTH-DZVP basis from the CP2K software package [132] is applied. MLWFs are constructed using Wannier90 [133], and interaction and hopping parameters are extracted via an in-house cRPA implementation [126].
For tPA, the initial DFT calculation is performed with 28 -points along the chain to ensure parameter convergence, while for PT, 12 -points are used. The direct interactions exhibit a slow convergence with respect to the number of -points, following a dependence (see Supporting Information). This behavior arises from the Coulomb divergence in the Gaussian density fitting scheme implemented in PySCF [134, 135]. As proposed by Berkelbach and coworkers [135], we extrapolated these parameters based on a reliable fit. The hopping and other interaction parameters do not suffer from this issue.
III.3 Tensor Network Calculations
Once the Hamiltonian is constructed, infinite two-site DMRG is performed to compute the ground state of the system while dynamically determining the bond dimension, followed by several VUMPS iterations to guarantee convergence. The relative tolerance is set to . Excitations on top of the ground state are determined using the quasiparticle algorithm, as introduced in Refs. [123, 124]. We look for the first particle and hole excitation as a function of the momentum and identify their sum as the required energy to create a free electron-hole pair. The band gap of the model is then extracted as the minimum of this energy over all momenta. In the case of an indirect band gap, the momenta of the particle and hole excitations may differ. However, this was not observed in the present work. Excitons, on the other hand, correspond to excitations in the trivial sector.
IV Results and Discussion
We present the results of our approach applied to tPA, PT, and Sr2CuO3. For tPA, a relatively simple system, we focus on the flexibility and quantitative accuracy that is offered by TNs. In the case of PT, we demonstrate the feasibility of effective models containing a large number of bands. Finally, we show that our approach is also applicable to quasi-1D materials such as Sr2CuO3.
IV.1 trans-Polyacetylene
Polyacetylene is the simplest conjugated polymer, existing in two configurations: trans- (see Figure 1a) and cis-polyacetylene. The cis isomer is thermodynamically less stable and converts to the trans form upon heating, allowing the preparation of pure trans samples for experimental studies [138]. Therefore, our discussion focuses on trans-polyacetylene.